OpenStudy (anonymous):

Which of the following is a solution to the equation log 6 x + log 6 (x – 1) = 1 ? Answer: A)x = 6/5 B)x = 3 C)x = 2 D)x = 4

OpenStudy (anonymous):

log6 x - log6 (x-1) =1 log6 (x/(x-1) = 1 6 power one = x/(x-1) 6 = x/(x-1) 6x - 6 = x 5x = 6 x = 6/5 = 1.2

OpenStudy (anonymous):

Thanks:)